### v1

# Pro-Pose: Unpaired Full-Body Portrait Synthesis via Canonical UV Maps

Abstract
Photographs of people taken by professional photographers typically present the person in beautiful lighting, with an interesting pose, and flattering quality. This is unlike common photos people can take of themselves. In this paper, we explore how to create a “professional” version of a person’s photograph, i.e., in a chosen pose, in a simple environment, with good lighting, and standard black top/bottom clothing. A key challenge is to preserve the person’s unique identity, face and body features while transforming the photo. If there would exist a large paired dataset of the same person photographed both “in the wild” and by a professional photographer, the problem would potentially be easier to solve. However, such data does not exist, especially for a large variety of identities. To that end, we propose two key insights: 1) Our method transforms the input photo and person’s face to a canonical UV space, which is further coupled with reposing methodology to model occlusions and novel view synthesis. Operating in UV space allows us to leverage existing unpaired datasets. 2) We personalize the output photo via multi image finetuning. Our approach yields high-quality, reposed portraits and achieves strong qualitative and quantitative performance on real-world imagery.
Figure 1: Pro-Pose Avatar Generation. From a single in-the-wild photo (top row), Pro-Pose generates a portfolio of high-fidelity avatars (bottom rows), fully driven by arbitrary SMPL-X poses [36]. The appearance is canonicalized to a black tank top and shorts and photoshoot setting environment. Our method preserves the user’s identity, facial features, and body shape across novel views and complex poses.
[…]
1 Introduction
Images have become the foundation of digital identity, yet everyday photos are often captured in uncontrolled conditions, with cluttered backgrounds, imperfect lighting, and arbitrary poses and clothing. Consequently, obtaining a clean, controllable “photoshoot-style” representation is difficult, primarily due to the scarcity of high-quality training data. Existing reposing methods such as MCLD [24] and LEFFA [53] rely on the small-scale paired DeepFashion dataset [25] ($\approx$100 identities), leading to overfitting and limited generalization to unseen individuals, and they cannot modify garments. In contrast, virtual try-on (VTO) models [9] typically rely on paired data of a source person and a target garment. Crucially, during training, they derive the person input from the target image and thus are trained to preserve the original pose rather than control or modify it. Consequently, neither method can jointly repose a person and generalize to real-world imagery - motivating our goal to generate photorealistic, reposed portraits of a person in standardized photoshoot conditions that preserve identity while abstracting away clothing complexity. As shown in Figure 1, our approach can produce a versatile portfolio of reposed, fashion-style portraits from a single input.
In this paper, we address the challenge of limited identity diversity by leveraging both abundant single images [16, 56, 55] and scarce paired data [25]. We propose a self-supervised framework utilizing a canonical UV-space [36], which theoretically decouples pose from texture. However, practical extraction remains ill-posed, since the specific boundaries of the visible texture map inevitably leak pose information. To eliminate these clues, we introduce Donor-based UV Reposing. This method uses occlusion masks from unrelated “donor” images to disguise the original visibility patterns, forcing the model to learn robust geometric warping. This formulation enables seamless joint training on 30K paired samples and $\approx$470K unpaired single images. Finally, we employ Gemini 2.5 Flash Image [7] to standardize the garments in the target images, creating a Base Clothing (BC) dataset featuring subjects in black tops and shorts with fixed faces and poses.
Our method employs a dual-branch training strategy to leverage both paired and unpaired data. The paired branch conditions the model on a SMPL-X [36] target, the partial UV texture map, and a source face crop to synthesize the subject in the standardized base garment. Conversely, for unpaired data, we utilize a self-supervised approach where the input also serves as the target. Here, we extract a donor-based UV map and drop out the face condition. This forces the network to reconstruct identity solely from the warped texture, preventing it from trivially copying input pixels. Finally, to address the overfitting risks inherent in single-shot personalization methods trained on limited data, we introduce a lightweight finetuning strategy. This adapts the pre-trained priors to new subjects using minimal reference images, ensuring high identity preservation without requiring large-scale multi-image datasets.
We summarize our contributions as follows: (1) a scalable joint training framework that integrates scarce paired data [25] with abundant unpaired single images [16, 56, 55] to overcome the limited identity diversity of existing benchmarks; (2) a novel self-supervised Donor-based UV Reposing mechanism that reduces coupling between pose and texture to prevent boundary-based pose leakage, enabling effective learning from unpaired data; (3) a stragety to create a large scale Base Clothing (BC[25, 16], synthesized via Gemini 2.5 Flash Image [7] to ensure consistent visual conditions; (4) a subject-specific adaptation mechanism that utilizes our unified training objective to fine-tune LORA layers at test-time, significantly reducing identity drift in challenging scenarios; and (5) extensive evaluations demonstrating that our approach achieves realistic, identity-faithful synthesis on challenging real-world imagery.
2 Related Work
2D Human Image Synthesis. Research in 2D human image synthesis has historically split into related tracks: pose-guided generation and disentangled attribute editing. Pose-guided methods aim to generate a novel pose of a person from a single image [41, 57, 21, 37]. To this aim, PG2 [29] adopted a coarse-to-fine pipeline, while subsequent single-stage architectures improved end-to-end training by incorporating spatial alignment via deformable skip connections [41], progressive attention [57], or dense appearance flow [21]. Parallel work focused on disentangling attributes, such as appearance and shape [5], or attribute-specific editing [31, 50]. This line of work was further generalized by motion models for arbitrary object animation [42]. To reduce paired data needs, [40] introduced cyclic self-supervision. However, these 2D models often struggle with spatial consistency and entangle pose with texture, leading to artifacts and identity loss.
3D-Aware Human Generation. To overcome spatial inconsistency, later methods incorporated 3D reasoning, implicit surface modeling, or neural rendering [38, 39, 12, 46, 52]. 3DHumanGAN [49] introduced a 3D-pose mapping module to enforce geometric plausibility, and EG3D [2] established efficient hybrid explicit/implicit 3D GANs for view-consistent generation. Other work such as PIFu [38] and PIFuHD [39] recovered pixel-aligned implicit fields, while ConTex-Human [8] focused on achieving texture consistency. More recent approaches have explored 3D Gaussian Splatting for real-time avatars [17, 32, 33] and self-supervised learning from social media videos [13, 14]. However, these approaches rely on strict multi-view or 3D supervision. While such datasets may be voluminous in terms of total frames, they remain severely limited in identity diversity compared to 2D images, preventing generalization to the distribution of real-world human appearances.
Diffusion-Based Animation. Recent diffusion-based techniques have significantly advanced fidelity and controllability [26, 24, 53, 11, 44, 28, 1]. Coarse-to-Fine Latent Diffusion (CFLD) [26] separates semantic and texture modeling; Animate Anyone [11] and DisCo [44] extend this paradigm to temporal consistency and disentangled control; and follow-up works such as MCLD [24] and LEFFA [53] enhance pose controllability. While high-quality, these models remain highly data-hungry, relying on large paired datasets like DeepFashion. Consequently, they are prone to overfitting and “identity drift” on unseen individuals.
Virtual Try-On and Garment Manipulation. Virtual try-on (VTO) methods focus on transferring garments onto target bodies while preserving identity [9, 43, 48, 19, 34, 47, 56, 55]. VITON [9] and CP-VTON [43] pioneered geometric matching for in-shop clothing. M&M VTO [56] extends this field with multi-garment diffusion control. However, these methods assume garment–person pairs rather than person-pose pairs, limiting their ability to handle reposing.
Summary. Existing 2D models lack consistency, 3D methods require expensive supervision, and diffusion animators rely on scarce paired data. To address these limitations, we propose a self-supervised framework formulated in UV space. By decoupling pose from texture and introducing donor-based reposing, we leverage abundant unpaired data alongside scarce paired data to generate photorealistic, reposed, and garment-neutral avatars from single images.
3 Method
We address the challenge of synthesizing highly realistic and pose-controllable human avatars by proposing a unified framework capable of learning from both abundant single-view images and scarce paired data. This integration is achieved through a novel supervision strategy formulated directly in canonical UV texture space, which naturally disentangles pose from appearance.
Figure 2: Pose Leakage Mitigation. (a) Standard partial textures $\mathbf{T_{p}}$ leak source pose information via occlusion boundaries, allowing trivial reconstruction shortcuts. (b) Generating pseudo-pairs via image-space rendering is computationally prohibitive for online training. (c) Our Donor-based Reposing efficiently bypasses rendering by applying random donor masks $\mathbf{M}_{\tilde{\mathbf{p}}}$ directly in UV space. This simulates novel occlusions, preventing leakage and forcing the network to learn robust geometric warping.
Figure 3: Overview of our Avatar Generation Framework. Our approach leverages single-view datasets by operating in a canonical UV space, extracting UV texture and pose [36]. Left (Paired Supervision): When ground-truth pose pairs are available, we condition the Flow Matching model directly on the target pose and face crop. Right (Single-View Self-Supervision): To prevent ”pose leakage” from occlusion boundaries when training on single images, we introduce a Donor-based UV Reposing module (Sec. 3.2.1). This synthetically re-poses the input texture using a random donor visibility mask, forcing the model to learn robust identity representations. Furthermore, we drop-out the face crop condition in this branch to prevent trivial reconstruction via pixel-perfect information leakage.
3.1 Overview and Problem Formulation
The fundamental goal is to generate an image, $\mathbf{A}_{\mathbf{p}}$, of a specific person in a novel target pose $\mathbf{p}$ (represented as a SMPL-X [36] rendering). Given a single reference image $\mathbf{I}_{\mathbf{p}^{\prime}}$ of that person in a source pose $\mathbf{p}^{\prime}$, we formulate this as a conditional generation problem:
$\mathbf{A}_{\mathbf{p}}=f_{\theta}(\mathbf{I}_{\mathbf{p}^{\prime}},\mathbf{p})$ (1)
where $f_{\theta}$ is the parameterized model we aim to learn.
A standard approach to train such a model involves minimizing a reconstruction loss, such as $\|\mathbf{A}_{\mathbf{p}}-\mathbf{I}_{\mathbf{p}}\|$, where $\mathbf{I}_{\mathbf{p}}$ is a ground-truth image of the same person in the target pose $\mathbf{p}$. This requires a large dataset of pose-paired images $(\mathbf{I}_{\mathbf{p}^{\prime}},\mathbf{I}_{\mathbf{p}})$. However, collecting such data at scale presents significant challenges due to scarcity, limited identity diversity, and inconsistencies (e.g., clothing changes) between pairs.
These limitations motivate leveraging abundant unpaired single images. However, naively setting $\mathbf{p}=\mathbf{p}^{\prime}$ leads to a trivial identity function. The central challenge is therefore to formulate a self-supervised manner that forces the model to learn meaningful representations from a single image.
3.2 Self-Supervision in Canonical UV Space
To enable effective learning from abundant single-view images, we shift our representation from image space to a canonical UV texture space. Ideally, we would extract a complete, identity-specific texture map $\mathbf{T}$ from any input image, perfectly decoupled from the subject’s pose. This would allow us to formulate a unified generator $g_{\theta}$ for any target pose $\mathbf{p}$:
$\mathbf{A}_{\mathbf{p}}=g_{\theta}(\mathbf{T},\mathbf{p})$ (2)
In practice, single-view unwrapping is ill-posed due to self-occlusions. We can only obtain a partial texture map, $\mathbf{T}_{\mathbf{p}}=\mathbf{T}\odot\mathbf{M}_{\mathbf{p}}$, where $\mathbf{M}_{\mathbf{p}}\in\{0,1\}^{H\times W}$ is the binary visibility mask uniquely defined by pose $\mathbf{p}$ (see Figure 2 a). We must therefore approximate the generator using a partial information:
$\mathbf{A}_{\mathbf{p}}\approx g_{\theta}(\mathbf{T}_{\mathbf{p}},\mathbf{p})$ (3)
Pose Leakage via Occlusion Boundaries.
While Eq. 3 enables self-supervision, it introduces a critical challenge: the partial texture $\mathbf{T}_{\mathbf{p}}$ is highly correlated with the pose $\mathbf{p}$ via the mask $\mathbf{M}_{\mathbf{p}}$. As visualized in Figure 2 a, occlusion boundaries in $\mathbf{T}_{\mathbf{p}}$ perfectly align with the target pose. This allows the network to minimize reconstruction loss trivially by ”copy-pasting” visible pixels based on these boundaries, rather than learning the desired 3D geometric warping.
3.2.1 Donor-Based UV Reposing
To mitigate pose leakage, we must break the correlation between the input texture boundaries and the target pose. While rendering the textured SMPL-X mesh into novel poses generates pseudo-pairs in image space (Figure 2 b), this process is computationally prohibitive for online training. We instead bypass explicit rendering by formulating the problem directly in UV space via Donor-based UV Reposing (Figure 2 c), an efficient 2D strategy that synthetically simulates novel pose occlusions.
[…]
$\mathbf{A}_{\mathbf{p}}\approx g_{\theta}(\mathbf{T}_{\mathbf{p}\rightarrow\tilde{\mathbf{p}}},\mathbf{p})$ (4)
3.3 Generative Framework
We parameterize our generator $g_{\theta}$ as a Latent Rectified Flow model [23], instantiated using the state-of-the-art Flux.1 [dev] transformer backbone [18]. Consistent with this architecture, we operate in the compressed 16-channel latent space defined by the pre-trained Flux Variational Autoencoder (VAE), comprising an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$.
Target images $\mathbf{I}_{\mathbf{p}}$ are compressed into latent representations $\mathbf{x}_{0}=\mathcal{E}(\mathbf{I}_{\mathbf{p}})$. Following the Flow Matching formulation [23], we model the generative process as an Ordinary Differential Equation (ODE) [3] that transports samples from a standard Gaussian noise distribution $\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to the data distribution $\mathbf{x}_{0}$.
We train a velocity prediction network $v_{\theta}$ to estimate the vector field driving this transport using the standard Flow Matching objective:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{c}}\left[\|v_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-(\mathbf{x}_{1}-\mathbf{x}_{0})\|^{2}\right]$ (5)
where $t\sim\mathcal{U}[0,1]$ is the timestep, $\mathbf{c}$ is the condition set, and $\mathbf{x}_{t}=t\cdot\mathbf{x}_{1}+(1-t)\cdot\mathbf{x}_{0}$ represents the linear interpolation between noise ($\mathbf{x}_{1}$) and data ($\mathbf{x}_{0}$).
Once trained, our high-level generator $g_{\theta}(\mathbf{c})$ (Eq. 4) is defined by numerically integrating this field backwards from noise ($t=1$) to data ($t=0$), followed by decoding:
$g_{\theta}(\mathbf{c})=\mathcal{D}\left(\mathbf{x}_{1}+\int_{1}^{0}v_{\theta}(\mathbf{x}_{t},t,\mathbf{c})dt\right)$ (6)
3.4 Training Data and Conditioning Strategy
We enable joint training via a unified condition vector $\mathbf{c}=\{\mathbf{T}_{\text{in}},\mathbf{p}_{\text{target}},\mathbf{I}^{FC}_{\text{in}}\}$, comprising the input texture, target pose, and an optional identity-anchoring face crop extracted via MediaPipe [27]. An overview of our unified training framework is presented in Figure 3. We instantiate $\mathbf{c}$ based on the data source:
3.4.1 Paired Supervision
For datasets with ground-truth pairs (reference $\mathbf{I}_{\mathbf{p}^{\prime}}$, target $\mathbf{I}_{\mathbf{p}}$), we condition on the reference partial texture and face crop to reconstruct the target view ($\mathbf{x}_{0}=\mathcal{E}(\mathbf{I}_{\mathbf{p}})$):
$\mathbf{c}_{\text{paired}}=\{\mathbf{T}_{\mathbf{p}^{\prime}},\mathbf{p},\mathbf{I}^{FC}_{\mathbf{p}^{\prime}}\}$ (7)
The face crop $\mathbf{I}^{FC}_{\mathbf{p}^{\prime}}$ is included to boost fidelity in facial regions where the partial texture $\mathbf{T}_{\mathbf{p}^{\prime}}$ may lack resolution.
3.4.2 Single-View Self-Supervision
For single-view images ($\mathbf{I}_{\mathbf{p}}$), we employ Donor-based Reposing (Eq. 4). We condition on the donor-masked texture $\mathbf{T}_{\mathbf{p}\rightarrow\tilde{\mathbf{p}}}$ and the target pose. Crucially, we explicitly drop the face crop condition ($\mathbf{I}^{FC}_{\text{in}}=\emptyset$) to prevent pixel-perfect information leakage:
$\mathbf{c}_{\text{single}}=\{\mathbf{T}_{\mathbf{p}\rightarrow\tilde{\mathbf{p}}},\mathbf{p},\emptyset\}$ (8)
This constraint forces the network to recover identity from the warped texture, ensuring robust geometric learning.
Figure 4: Finetuning pipeline. We build input–target pairs from a few-shot subject set and apply a facially masked paired Flow Matching loss to personalize the model at test time.
3.5 Test-Time Personalization via Few-Shot Adaptation
While our method already achieves high identity preservation in a feed-forward manner, we can further enhance fidelity through test-time personalization via few-shot adaptation. Given a small set $\mathcal{S}$ comprising $N$ images of a person in various poses, we create pairs by using one view as the reference and another as the target. We then fine-tune our LoRA adapters using the paired objective, as illustrated in Figure 4. This personalizes the model to the specific identity, improving consistency, particularly for extreme pose generation.
We constrain this personalization to the face by applying the loss exclusively to the facial region using a binary mask $\mathbf{M}_{j}$, which represents the target’s latent-space face mask:
$\mathcal{L}_{\text{FT}}=\mathbb{E}_{t,(i,j)}\left[\|\mathbf{M}_{j}\odot\left(v_{\theta}(\mathbf{x}_{j,t},t,\mathbf{c}_{\text{paired}})-v_{\text{target}}\right)\|^{2}\right]$
where $\mathbf{c}_{\text{paired}}=\{\mathbf{T}_{i},\mathbf{p}_{j},\mathbf{I}^{FC}_{i}\}$ is the conditions extracted from a sample pair $(i,j)$ drawn from $\mathcal{S}$ (such that $i\neq j$), and $v_{\text{target}}=\mathbf{x}_{j,1}-\mathbf{x}_{j,0}$ is the target transport guidance.
3.6 Implementation Details
Figure 5: Base Clothing (BC) Standardization. We apply different preprocessing strategies based on the data. For DeepFashion [25] and Commerce images [56, 55], we generate the base garment while preserving pose and identity via pixel-aligned editing (Prompt V1). For FFHQ [16], we use generative outpainting to expand limited face crops into full-body samples (Prompt V2).
Basic Clothing (BC) Dataset. We construct a standardized training dataset by processing paired DeepFashion [25] data, unpaired FFHQ [16] faces, and commercial images [56, 55] using Gemini 2.5 Flash Image [7]. As shown in Figure 5, we employ distinct strategies: pixel-aligned editing for full-body inputs to standardize garments to a black tank top and shorts, and generative outpainting to expand face crops into full-body portraits. The specific text prompts are detailed in the Appendix.
DeepFashion (In-Domain)
[…]
CFLD [26] 17.645 7.149 0.7478 0.1819 0.48 0.3180 0.9731 4.15
MCLD [24] 18.211 7.079 0.7557 0.1757 0.49 0.3440 0.9654 4.29
LEFFA [53] 14.020 4.23 0.7550 0.1190 0.44 0.5794 0.9409 4.41
OnePoseTrans [6] 13.570 8.740 0.6053 0.3066 0.46 0.5750 0.9476 4.32
UniHuman [20] 14.050 6.250 0.7960 0.1560 0.46 0.5810 0.9434 4.03
Unpaired Only 15.560 7.780 0.7013 0.2455 0.47 0.3400 0.9189 4.23
Paired Only 19.350 4.23 0.8094 0.0760 0.48 0.6103 0.9757 7.25
Ours (Unpaired + paired) 19.330 4.24 0.8095 0.0780 0.48 0.5919 0.9712 7.23
WPose (Out-of-Domain)
[…]
CFLD [26] 15.43 96.07 0.744 0.208 0.31 0.0885 0.6412 1.94
MCLD [24] 15.64 94.23 0.759 0.201 0.35 0.0995 0.6478 1.96
LEFFA [53] 16.71 67.85 0.776 0.193 0.32 0.0914 0.5725 2.01
OnePoseTrans [6] 17.23 27.43 0.818 0.151 0.33 0.1735 0.7205 4.44
UniHuman [20] 17.64 27.748 0.807 0.161 0.34 0.1121 0.7207 2.89
Unpaired Only 16.06 6.98 0.7565 0.2179 0.37 0.3800 0.6748 4.26
Paired Only 18.23 6.73 0.8143 0.1615 0.34 0.4954 0.6915 7.23
Ours (Unpaired + paired) 19.87 6.04 0.849 0.127 0.38 0.5563 0.7383 7.49
Table 1: Quantitative evaluation and ablation study. We report results on DeepFashion (top) and WPose (bottom). The upper section of each block compares our method against state-of-the-art baselines, while the lower section details our ablation study (Unpaired Only, Paired Only, and our full hybrid model). Red indicates the best performance and Blue indicates the second best across all methods.
Training Setup. Our combined dataset totals approximately 500k samples: 470K single unpaired images [16, 56, 55], and 30K paired images from DeepFashion dataset[25] (We restrict training to a subset of DeepFashion data to prevent overfitting to its limited number of unique identities ($\approx 100$), while ensuring all identities are still represented.)
Optimization. We use Flux.1 [dev] [18] as our backbone, training rank-128 LoRA adapters using AdamW with a learning rate of $10^{-4}$. Training proceeds for 100K iterations on 128 TPUv5 chips with a global batch size of 128.
Donor Pool & Dropout. For every single-view sample, we pre-compute a pool of 10 valid donor masks $\mathbf{M}_{\tilde{\mathbf{p}}}$ that satisfy an IoU constraint of $[0.4,0.8]$ with the source mask. To prevent modality over-reliance, we apply mutually exclusive conditioning dropout: we either drop all conditions ($p=0.05$), or individual components: $\mathbf{T}_{\text{in}}$ ($p=0.3$), $\mathbf{I}^{FC}_{\text{in}}$ ($p=0.3$), or $\mathbf{p}_{\text{target}}$ ($p=0.1$).
4 Experiments
In this section, we lay out the details of our evaluation strategy, including datasets, baselines, and metrics used to compare our method against state-of-the-art approaches. We also conduct extensive ablations to highlight the effectiveness of each of our contributions.
Method PSNR FID SSIM FaceSim DINO HPSv3
Base 18.59 6.88 0.823 0.4837 0.689 7.15
Finetuned 19.34 6.76 0.835 0.5722 0.705 7.17
Table 2: Ablation study (Finetuning). Finetuning can lead to better face and identity preservation compared to our base model.
4.1 Datasets and Metrics for Evaluation
Following prior work [24, 26, 53], we use the DeepFashion In-Shop Clothes Retrieval test split [25] (8570 image pairs). To assess generalization to in-the-wild scenarios, we additionally evaluate on WPose dataset [20] (2305 image pairs).
Since our method canonicalizes clothing, direct comparison against ground-truth images (which include diverse garments and backgrounds) is not meaningful. For image similarity (PSNR, SSIM [45]) and perceptual metrics (FID [10], LPIPS [51]), we therefore compare our outputs against a Basic Clothing version. We further report these metrics on the original image only face regions in the appendix.
Pose, Identity, Semantic and Perceptual Metrics. To measure how well the generated avatar matches the target pose, we compute Object Keypoint Similarity (OKS) [22] between the predicted and ground-truth keypoints. Identity fidelity is evaluated within the facial region. We compute Face Similarity (FaceSim) using cosine similarity of ArcFace [4] identity embeddings. We report DINOv2 similarity [35] as a measure of semantic and structural alignment. For reference-free perceptual quality, we report HPSv3 [30].
Figure 6: Ablation Study of Training Data Sources. We evaluate the impact of different training dataset combinations on generated avatar quality. The result columns display models trained on Unpaired data only, Paired data only, and the full Unpaired + Paired combination, respectively. One can observe the improvement in fidelity with the full combined dataset.
Figure 7: Finetuning Qualitative results. Although the input image suffered from motion blur, leading to blurry base model results, finetuning successfully restored crisp, detailed face features, demonstrating the model’s ability to learn robust features from multiple images rather than relying solely on one input.
4.2 Quantitative and Qualitative Comparison
We compare our method against the recent diffusion-based pose-conditioned avatar generation models, including MCLD [24], CFLD [26], LEFFA [53], OnePoseTrans [6], and UniHuman [20].
As detailed in Table 1, our method achieves SOTA performance on both benchmarks. On the in-domain DeepFashion set [25], our model sets a new SOTA, significantly outperforming all baselines across image fidelity (PSNR, SSIM, LPIPS), identity (FaceSim), and perceptual quality (HPSv3). For the in-the-wild WPose dataset [20], where we use foreground-masked metrics (M-PSNR, M-SSIM, M-LPIPS), our full model (Unpaired + Paired) conclusively dominates all SOTA baselines. This highlights the critical role of our unpaired data strategy in achieving robustness for challenging, in-the-wild data.
Figure 9 provides a qualitative comparison with SOTA methods. As demonstrated, our method generates high-fidelity avatars that accurately preserve the input person’s identity and facial features. For instance, the eighth row shows robust preservation of face and beard shape. Furthermore, our approach more robustly follows the target pose than any other method (e.g., sixth row) and maintains body shape fidelity relative to the original image (e.g., ninth row).
4.3 Ablation Study
Hybrid Data Strategy. We validate our hybrid data strategy by training three variants of our model: (i) Unpaired Only, trained exclusively on unpaired single images; (ii) Paired Only, trained exclusively on paired images; and (iii) Unpaired + Paired, our full model.
As shown in Table 1, while the Paired Only model performs well on in-domain data (DeepFashion [25]), it suffers on out-of-domain the WPose set [20], confirming its poor generalization. Figure 6 demonstrates that training solely on the limited paired dataset leads to overfitting and identity leaking from the training data. Alternatively, the model trained exclusively on unpaired data fails when subjected to significant pose changes. Our full method (Unpaired + Paired) synergizes these sources, effectively utilizing the larger dataset to learn robust identity priors while maintaining stable pose guidance.
Test-Time Personalization. Furthermore, we validate our test-time personalization strategy, detailed in Section 3.5. As shown in Table 2 and qualitatively in Figure 7, applying this few-shot adaptation further improves performance over our base model across all metrics. The most significant gain is in identity preservation, with FaceSim improving by 18.3%. This confirms that our finetuning approach is highly effective at enhancing model personalization and identity fidelity for specific subjects. We share more finetuning results in the Appendix.
Figure 8: Limitations of the base model. Left: Extreme pose changes (e.g., back-to-front) result in minimal UV texture overlap, forcing the model to hallucinate unseen regions. While general attributes like beard shape or hair color may be inferred from partial cues, precise facial structure can lose fidelity. Right: Because our explicit UV conditioning is restricted to the face, body-specific skin details such as tattoos are not preserved.
Figure 9: Qualitative Comparison. The first two rows display results from the DeepFashion dataset [25], and the subsequent rows utilize real images from the WPose dataset [20]. The columns show the Original Person, Target Pose, and results from state-of-the-art methods (MCLD [24], Leffa [53], UniHuman [20], OnePoseTrans [6]) compared to Ours and the Ground Truth (GT). Our method demonstrates superior performance in preserving the person’s identity, particularly the facial characteristics, across varying poses. For instance, in the third and fifth rows, previous methods introduce significant identity distortion and loss of facial likeness (e.g., changes in jawline or features). In contrast, Ours robustly preserves the unique face structure and identity of the original person, closely matching the GT result.
5 Limitations and Future Work
While our framework significantly advances single-image avatar reposing, some limitations remain. First, the base model can struggle with extreme pose changes (Figure 8, left), where minimal texture overlap forces identity-reducing hallucination. Although multi-image fine-tuning mitigates this issue, the base model may still fail in challenging cases. Second, our UV conditioning is restricted to the face; thus, distinctive body features like tattoos (Figure 8, right) or birthmarks cannot be represented on generic skin regions. Future work will address these limitations by extending UV conditioning to full-body skin texture and applying the skin- preservation during finetuning stage.
6 Conclusion
In this paper, we presented a novel framework for generating photorealistic, reposed human avatars from unconstrained single images. By introducing a self-supervised Donor-based UV Reposing strategy, we effectively decoupled pose from texture, enabling our Flow Matching generator to learn robust identity preservation from massive unpaired image collections. We further showed that this strong base model can be rapidly specialized via multi-image fine-tuning to handle more extreme reposing scenarios. By standardizing data into a base clothing and formulating generation in canonical UV space, our approach provides a scalable foundation for creating high-fidelity, controllable digital avatars from everyday photography.
[…]
Supplementary Material
Figure 10: Pro-Pose In-the-wild. We apply our method to real-world images characterized by varying poses and environments. The columns display (from left to right): (1) the original input image; (2) the target pose represented by a SMPL-X mesh (reposed using ground truth parameters); (3) the output generated by Pro-Pose with a corresponding face crop for detailed visual comparison. Our method consistently preserves facial details and structure compared to the ground truth.
7 Base Clothing (BC) Dataset.
We utilize paired data from DeepFashion [25], unpaired single images from FFHQ [16], and commercial data derived from e-commerce websites [56, 55]. To ensure visual uniformity across these diverse sources, we generate the Base Clothing (BC) dataset using an identity-preserving I2I model (Gemini 2.5 Flash Image [7]), as shown in Figure 5 (main paper). Depending on the nature of the input data, we employ either Prompt V1 or Prompt V2:
Prompt V1:
Precisely edit IMAGE_0 to change the garment worn by the person in the photo to a black sleeveless tank top and black shorts. Do not change the person in the image or the framing. Remove or replace any accessories such as bags, hats, jewelry and sunglasses in IMAGE_0. Do not alter IMAGE_0 in any other way. Keep it pixel-aligned.
Prompt V2:
Precisely edit IMAGE_0 to generate a full body image of this person (head to toes) while keeping the face and head pose same in a black sleeveless tank top and 3-inch black shorts. Generate the body pose suiting the existing face pose. The person is posing naturally against a clean and plain background. Even, soft studio lighting illuminates the subject from the front, highlighting the form and texture of the garments. No accessories (no bags, hats, jewelry, sunglasses). Professional product photography style, no harsh shadows or distracting elements. High resolution, sharp focus.
8 Canonical UV-Space Formulation
To formulate our generative objective in the canonical UV space, we first extract the parametric body model and the corresponding partial texture map $\mathbf{T}_{\mathbf{p}}$ from the input image $\mathbf{I}$.
[…]
Figure 11: Random Patch Masking Augmentation vs. Ours (Donor-Based UV Reposing). We compare random patch masking texture augmentation against our donor-based UV reposing method. With standard random masking, the shape of the UV boundaries inadvertently reveals pose information. This leakage causes the model to generate significant artifacts, particularly near face boundaries. As seen in the first row, hand-face occlusions create hand artifacts, while the second and third rows display hallucinated accessories such as swimming goggles, hats, and dreadlocks. By effectively decoupling texture from pose via donor-based UV reposing, our method resolves these issues, yielding cleaner results with superior identity preservation.
Partial Texture Extraction. To obtain the partial texture $\mathbf{T}_{p}$, we perform an inverse rendering operation via barycentric sampling in UV space. Let $\mathbf{V}^{uv}\in\mathbb{R}^{N\times 2}$ denote the fixed UV coordinates of the SMPL-X topology, and $\mathbf{V}^{img}\in\mathbb{R}^{N\times 2}$ denote the projection of the 3D vertices $\mathbf{V}$ onto the coordinate space of input image $\mathbf{I}$.
Computing $\mathbf{T}_{\mathbf{p}}$ directly via ray-casting is computationally expensive. Instead, we adopt an efficient rasterization-based approach. We rasterize the mesh in UV space, utilizing $\mathbf{V}^{uv}$ as the position attributes. During rasterization, we interpolate the image-space coordinates $\mathbf{V}^{img}$ across the triangle faces. For a given texel $\mathbf{u}$ in the canonical UV space $\Omega_{uv}$, located within a triangle $f\in\mathcal{F}$ with barycentric weights $\mathbf{b}=(b_{1},b_{2},b_{3})$, the corresponding source image coordinate $\mathbf{x}_{src}$ is computed as:
$\mathbf{x}_{src}(\mathbf{u})=\sum_{k=1}^{3}b_{k}\cdot\mathbf{V}^{img}_{f,k}$ (10)
where $\mathbf{V}^{img}_{f,k}$ corresponds to the image-space coordinate of the $k$-th vertex of face $f$. The pixel value for the partial texture at location $\mathbf{u}$ is then obtained by bilinear sampling of the input image:
$\mathbf{T}_{p}(\mathbf{u})=\begin{cases}\mathbf{I}(\mathbf{x}_{src}(\mathbf{u}))&\text{if }\mathbf{M}_{\mathbf{p}}(\mathbf{u})=1\\ 0&\text{otherwise}\end{cases}$ (11)
Here, $\mathbf{M}_{\mathbf{p}}\in\{0,1\}^{H\times W}$ is the binary visibility mask determined by the z-buffer during the initial estimation. To ensure high-fidelity unwrapping and avoid artifacts at UV seams, we filter triangles where the edge lengths in UV space exceed a threshold $\tau_{dist}$.
9 Ablation Study: Donor-based UV Reposing vs Random Patch Masking
In Section 3.2.1 (main paper), we introduced Donor-based UV Reposing to prevent the model from learning shortcuts. We argued that using a real mask from a ”donor” creates realistic occlusions that force the network to learn geometry. In this section, we test if this is actually necessary. We compare our method against a simpler baseline where we just block out random parts of the texture to create a synthetic pair.
Implementation Details. For the ”Random masking” baseline, we take the input facial texture ($\mathbf{T}_{\mathbf{p}}$) and cover it with random black squares. We use up to 6 non-overlapping patches, each $64\times 64$ pixels. All other training settings remain identical to our main method.
Qualitative Results. Visual comparisons in Figure 11 show that random masking is insufficient. The problem is that random squares do not change the overall ”shape” of the visible texture. The boundaries of the valid texture still look like the source pose. Because the model can still recognize the source pose from these boundaries, it tries to find a shortcut. However, the random black squares create confusing holes that don’t match natural occlusions. To explain these unnatural missing regions, the model, more frequently, hallucinates accessories.
Our Donor-based method avoids this because the donor mask looks like a real human pose. It changes the texture boundaries completely, so the model cannot rely on the visible pose outline and must learn beyond simple ”copy-pasting”.
Quantitative Results. Table 3 confirms these visual findings. The Random Masking baseline performs worse on all metrics. Most notably, identity preservation drops (FaceSim decreases by 6.75%) and overall image quality degrades. This proves that simply hiding pixels is not enough; the occlusion must mimic a realistic pose to train the model effectively.
[…]
Random masking 18.56 6.45 0.821 0.5187 0.6965 7.35
Ours 19.87 6.04 0.849 0.5563 0.7383 7.49
Table 3: Donor Masking Ablation. We compare our method against a ”Random Masking” baseline, that applies multiple black patches ($64\times 64$) to the input texture. The superior performance of our method confirms that structurally semantic donor masks are essential for learning robust geometry, whereas random patches fail to prevent boundary-based pose leakage.
10 Additional Results
Figure 12: Qualitative Comparison. The last three rows display results from the DeepFashion dataset [25], and all previous rows utilize real images from the WPose dataset [20]. The columns show the Original Person, Target Pose, and results from state-of-the-art methods (MCLD [24], Leffa [53], UniHuman [20], OnePoseTrans [6]) compared to Ours and the Ground Truth (GT). Our method demonstrates superior performance in preserving the person’s identity, particularly the facial characteristics, across varying poses. For instance, in the third and fifth rows, previous methods introduce significant identity distortion and loss of facial likeness (e.g., changes in jawline or features). In contrast, Ours robustly preserves the unique face structure and identity of the original person, closely matching the GT result.
Figure 13: Ablation Study of Training Data Sources. We evaluate the impact of different training dataset combinations on generated avatar quality. The result columns display models trained on Unpaired data only, Paired data only, and the full Unpaired + Paired combination, respectively. One can observe the improvement in fidelity with the full combined dataset.
DeepFashion (In-Domain)
[…]
CFLD [26] 14.34 0.7012 0.1823 0.9543
MCLD [24] 15.76 0.7151 0.1798 0.9429
LEFFA [53] 13.95 0.6956 0.12 0.9316
OnePoseTrans [6] 13.12 0.5738 0.322 0.9319
UniHuman [20] 13.87 0.7557 0.161 0.9534
Unpaired Only 12.87 0.6766 0.2489 0.9001
Paired Only 16.45 0.7765 0.087 0.9555
Ours (Unpaired + paired) 16.38 0.7785 0.089 0.9554
WPose (Out-of-Domain)
[…]
CFLD [26] 15.12 0.741 0.279 0.6109
MCLD [24] 15.29 0.757 0.24 0.6108
LEFFA [53] 16.47 0.77 0.236 0.5541
OnePoseTrans [6] 17.1 0.807 0.174 0.6719
UniHuman [20] 17.49 0.802 0.179 0.6814
Unpaired Only 15.81 0.748 0.255 0.6445
Paired Only 17.98 0.7866 0.165 0.6443
Ours (Unpaired + paired) 19.57 0.832 0.138 0.7043
Table 4: Quantitative Evaluation and Ablation Study. We compute PSNR, SSIM, LPIPS and DINO only on the face region of the generated images.
10.1 Extended Evaluation In-The-Wild
To demonstrate the robustness of our approach, we evaluate Pro-Pose on “in-the-wild” images sourced from CelebA-HQ [15]. These samples feature diverse lighting conditions, backgrounds, and subjects. The qualitative results, presented in Figure 10, confirm that our method generalizes effectively to unseen real-world data, maintaining high identity fidelity and pose accuracy.
10.2 Extended Qualitative Comparisons
We provide more extensive qualitative comparisons against state-of-the-art methods (MCLD [24], Leffa [53], UniHuman [20] , OnePoseTrans [6]) in Figure 12. Our method consistently produces higher fidelity faces and fewer geometric artifacts than competing approaches, particularly for challenging poses and diverse identities. Our method performs equally well on both in-domain (DeepFashion [25]) and in-the-wild (WPose [20]) scenarios, maintaining consistent quality across datasets.
10.3 Extended Quantitative Evaluation
To further validate the fidelity of our generated avatars, we report metrics computed exclusively on the facial regions in Table 4. This localized evaluation prevents the large, uniform areas of background and body from inflating the pixel-level metrics, thereby providing a stricter assessment of identity and facial detail preservation.
Metric Selection. We focus on PSNR, SSIM, LPIPS, and DINO to assess reconstruction quality and semantic alignment. We exclude OKS from this specific analysis, as it relies on full-body keypoints and becomes unstable when restricted to the sparse keypoints available in tight facial crops. Similarly, we omit HPSv3, as it is designed to assess global image aesthetics; when applied to small facial crops, it lacks the necessary global context to yield meaningful discriminative trends. Note that FaceSim is excluded from this table as it was already reported in the main paper (where it is inherently computed on facial embeddings).
Figure 14: Personalized Pro-Pose. We fine-tune our method using the images shown in the leftmost column (rows 1-4). Column (5) displays the target pose represented by a SMPL-X mesh (reposed using ground truth parameters). Columns (6) and (7) present the generated output without and with fine-tuning, respectively. The subsequent columns repeat this arrangement for a second target pose.
Analysis. The results reinforce the generalization gap observed in our main experiments. On the in-domain DeepFashion dataset, the ”Paired Only” model performs marginally better, consistent with its tendency to overfit to the training distribution. However, on the challenging, out-of-distribution WPose dataset, Pro-Pose significantly outperforms all baselines and the ”Paired Only” variant across all metrics. This confirms that our hybrid training strategy yields superior generalization on fine-grained facial details, avoiding the identity degradation observed in baselines.
10.4 Impact of Training Data Source
Complementing the quantitative ablation in the main paper, Figure 13 visually demonstrates the impact of our hybrid data strategy. Models trained on ”Unpaired Only” data struggle with extreme poses, while ”Paired Only” models suffer from identity drift. Our combined approach (”Unpaired + Paired”) leverages the strengths of both, achieving robust pose control and identity preservation.
10.5 Personalization with Few-Shot Learning
As outlined in Section 3.5 of the main paper, we perform few-shot adaptation to specialize the model for a specific target identity. Here, we provide the specific implementation details for this process.
Data Source and Preprocessing. While the visualization (Figure 4) in the main paper utilizes frames from a video sequence, our fine-tuning protocol is data-agnostic. It requires only a set of paired images of the subject in different poses; these can be sourced from a video or a collection of independent multi-view photographs. Crucially, before fine-tuning, we process all reference images to match our Base Clothing (BC) standard (black tank top and shorts). This ensures that the input data aligns with the canonical distribution the model was trained on, allowing the optimization to focus on identity features rather than clothing discrepancies.
Implementation Details. We fine-tune the pre-trained LoRA adapters for 5,000 iterations with a batch size of 4. To prevent the model from overfitting to the limited few-shot examples or forgetting its general geometric priors, we include a small set of regularization images in the training batches. For consistency and reproducibility, we maintain all other hyperparameters (including learning rate and optimizer settings) identical to the main training phase. While subject-specific hyperparameter tuning could theoretically yield higher fidelity results, we report all metrics using this fixed configuration.
More Qualitative Evaluation. We demonstrate how our few-shot learning strategy generalizes to in-the-wild scenarios [54] in Figure 14. Notably, the identities of the generated new poses become significantly closer to the original subject after fine-tuning, validating the effectiveness of this adaptation step.
Figure 15: DeepFashion. We highlight the top 100 identity clusters found in the DeepFashion dataset. We also exclusively highlight the first person on the top-left, who can be repeatedly obersved in the generations of the overfitted models.
Figure 16: Diagnosing Identity Overfitting via Cross-Dataset Generation. We evaluate the extent of identity overfitting by comparing the face similarity between images generated on the out-of-distribution WPose dataset and the identities present in the DeepFashion training set. Two Pro-Pose models are tested: Pro-Pose (paired only), trained on limited DeepFashion paired data, and Pro-Pose (unpaired + paired), trained on abundant unpaired data plus the limited paired data. The Pro-Pose (paired only) model exhibits a significantly higher face similarity to the DeepFashion training identities when reposing WPose identities. This result acts as a diagnostic, demonstrating that the paired-only model has overfit to the limited DeepFashion identities, causing it to generate new faces that are biased towards its training set, whereas the combined training approach ensures better generalization.
11 Analysis of Identity Overfitting
A primary motivation for our work is the limited identity diversity in existing paired datasets like DeepFashion. As visualized in Figure 15, DeepFashion contains limited ($\approx 100$) unique identities. Training generative models exclusively on such limited data leads to severe overfitting, where the model biases generated faces toward the training identities rather than the input subject.
11.1 DeepFashion Identity Analysis
To empirically validate the limited identity diversity of the DeepFashion In-Shop Clothes Retrieval benchmark [25], we performed a graph-based clustering analysis on the training set. Our pipeline proceeds as follows:
  1. 1.
    Embedding Extraction: We utilized ArcFace [4] to detect faces and extract identity embeddings. To ensure embedding reliability, we explicitly filtered out faces with extreme head poses (yaw or pitch $45^{\circ}$).
  2. 2.
    Graph Construction: We constructed an undirected graph where nodes represent images. Edges were established via a dual-criteria strategy:
    • Visual Similarity: An edge is created if the cosine similarity between two embeddings exceeds a strict threshold of $0.6$.
    • Structural Verification: We leveraged the dataset’s directory structure (where images in the same garment folder typically depict the same subject) to add ground-truth edges. However, to filter labeling errors or occlusions, these structural edges were only added if the similarity exceeded a verification threshold of $0.4$.
  3. 3.
    Clustering: We computed the connected components of this graph. This process yielded approximately 100 distinct clusters (identities), confirming that the dataset is heavily dominated by a small number of professional models appearing across thousands of garment items.
11.2 Identity Analysis of Generated Images
We diagnose this phenomenon in Figure 16. We reposed subjects from the out-of-distribution - WPose dataset using two models: one trained only on paired DeepFashion data, and our full model trained on both paired and unpaired data. We then measured the face similarity between the generated outputs and the DeepFashion training set.
  • The Paired-Only model (red histograms in Figure 16 (a) and (b)) produces outputs with high similarity to the training data, indicating it is ”pulling” novel identities toward the few identities it memorized during training.
  • Our Full Model (blue histograms in Figure 16 (a) and (b)) maintains lower similarity to the training set, proving that the inclusion of abundant unpaired data enables better generalization to unseen identities.
The box-plot in Figure 16 (c) further highlights the extensive number of samples generated by the paired-only model that have very high similarity scores to DeepFashion training samples.
The cumulative distribution (Figure 16 (d)) highlights that our model trained on both paired and unpaired data generated samples with a maximum Face Similarity score of $0.4128$, while almost $15\%$ of the total samples generated by the paired only model scored an even higher face similarity score.
11.3 DeepFashion Identity Limitation Leading to Overfitting
We highlight a dominant identity from the DeepFashion training set in the red box of Figure 15. Comparing this to the ”Paired Only” results in Figure 13 (and Figure 6 of main paper) reveals severe overfitting: As we can see in some cases the model tends to collapse to this specific training identity (for example second row in Figure 13). This bias is so strong that it overrides semantic gender cues, frequently substituting female subjects with this specific male face, confirming that the paired-only model relies on memorization rather than generalization.